OpenStudy (calebb8133):

i HAVE TO get this right The leg of a right triangle is 2 units and the hypotenuse is 3 units. What is the length, in units, of the other leg of the triangle? Square root of 5 units Square root of 13 units 5 units 13 units

OpenStudy (calebb8133):

@steve816

OpenStudy (calebb8133):

@steve816

OpenStudy (calebb8133):

HELP

OpenStudy (calebb8133):

@rishavraj

OpenStudy (calebb8133):

i think it is C

OpenStudy (calebb8133):

i am just double checking my answers

rishavraj (rishavraj):

Use Pythogores theorem ..and its not C

OpenStudy (calebb8133):

:(

OpenStudy (calebb8133):

thats all the help i get @rishavraj

OpenStudy (calebb8133):

...

rishavraj (rishavraj):

see  h = 3 y = 2 \[x = \sqrt{3^2 - 2^2}\]

OpenStudy (tommynaut):

a^2 + b^2 = c^2 Here, c = 3 (the hypotenuse) and a = 2 (we could have b = 2 instead, it doesn't make a difference) so then we have 2^2 + b^2 = 3^2 2^2 = 4, 3^2 = 9, so... 4 + b^2 = 9 b^2 = 9-4 Maybe you can figure out the rest

OpenStudy (vijeya3):

LOL it's square root of 5 that is the first option.... @rishavraj bechare bacho ke saath itna bhi strict mat ban bex'D

rishavraj (rishavraj):

@Vijeya3 lol dhyan rakhunga :P

OpenStudy (calebb8133):

ok thx everyone

OpenStudy (vijeya3):

;) good sweety

OpenStudy (calebb8133):

is he right @rishavraj

OpenStudy (calebb8133):

i have to get this right thats y i am double checking

OpenStudy (vijeya3):

lol dekh wo kitna trust krta hai tujhex'd

rishavraj (rishavraj):

its A ...go on @calebb8133 :)

OpenStudy (calebb8133):

thx :)

OpenStudy (calebb8133):

i may be back :)